# Universality in the velocity jump in the crack propagation observed for food-wrapping films for daily use

### v1

@@ Abstract
The velocity jump found in the crack propagation for rubbers has been a powerful tool for developing tough rubber materials. Although it is suggested by a theory that the jump could be observed widely for viscoelastic materials, the report on a clear jump is very limited and, even in such a case, reproducibility is low, except for elastomers. Here, we use a mundane food-wrapping film as a sample and observe the crack propagation velocity with pulling the sample at a constant speed in the direction perpendicular to the crack. As a result, we find the jump occurs at a critical strain with high reproducibility. Remarkably, the plot of the crack-propagation velocity as a function of strain can be collapsed onto a master curve by an appropriate rescaling, where the master curve is found to be universal for change in the pulling speed and in the sample height. The result reveals a key parameter for the jump is the strain, suggesting the existence of a small length that governs the deformation along the crack. The present study sets limitations on future theories and opens an avenue for the velocity jump to become a tool for developing a wide variety of tough polymer-based materials.
Introduction — To develop tough polymer-based materials are crucial in our modern life. Accordingly, the fracture of polymer based materials has been studied intensively [1, 2, 3, 4, 5, 6], but remains an active area of research. One of the principal topics in the study of polymer fracture is the crack propagation, which can be studied through various methods: the cyclic tensile test [7, 8], trouser test [9, 10], static tensile test [11, 12], and so forth [13, 14].
@@
In 2016, a theoretical explanation was proposed based on a simple viscoelastic model, which suggests the jump can be observed for a wide range of viscoelastic materials [18]. The jump was reproduced in numerical simulation [17], and the physical picture for the origin of the velocity jump proposed in Ref[18], the glass transition at the crack tip, is strengthened by a study performed under a collaboration among theoretical, numerical and experimental groups [20].
@@
As a matter of fact, the velocity jump was not observed for the porous polypropylene, when the crack propagation was observed under the static boundary condition [24]. However, the jump was observed in a dynamic experiment, in which the crack propagation was observed under a dynamic boundary condition [23]: we first introduce a short crack at one edge of the sample, and then start stretch the sample at a fixed speed to observe the crack propagation [25]. The reason the jump was not observed under the static experiment may be that the sample is relaxed during the preparation for giving a fixed stretch, while such effects of relaxation can be reduced in the dynamic experiment. This is one of the advantages of the dynamic experiment over the static one.
@@
Here, we use a mundane food-wrapping film as a sample and observe the crack propagation under the dynamic boundary condition. As a result, we find the jump occurs at a critical strain with high reproducibility. The crack-tip position as a function of time is obtained using two cameras, one for capturing the slow crack propagation before jump and the other with high frame rate for the fast crack propagation after jump. The plot of the crack-propagation velocity as a function of strain thus obtained shows a little sample dependence: for a fixed pulling velocity and a fixed sample size, we obtained quantitatively similar curves for different samples. Furthermore, we find that the velocity curve is independent from the pulling velocity but dependent on the sample height, while they can be collapsed onto a master curve by a rescaling: the master curve thus obtained is universal for change in the pulling speed and the height of the sample. The result reveals a key parameter for the jump is the strain, suggesting the existence of a small length that governs the deformation along the crack. The present study sets limitations on future theories and opens an avenue for the velocity jump to become a tool for developing a wide variety of tough polymer-based materials.
@@
From the strain-stress curve obtained at the room temperature at the strain rate $\simeq 0.02$ to 0.07 s-1, the elastic modulus is around 500 MPa and increases with the strain rate. From rheological measurements, the sample exhibits a glass transition at the time scale around tens of milliseconds at the room temperature, where the elastic modulus at the glass phase $\simeq 5$ GPa drops to a smaller value. See Appendix for further details on mechanical and rheological properties.
@@
Experimental — As illustrated in Fig. 1, we cut a sheet sample from a role of width $W$ (=150=150 mm) and set the sheet bottom to the fixed clamp and the sheet top to the mobile clamp so that the sample height at the zero strain becomes $L$ (=10=10 to 25 mm), i.e., the distance between the bottom and top clamps is set to $L$ at the zero strain. We introduce a short line crack of length $a$ at one of the free edge, before we start stretching the mobile clamp at a constant speed $U$ (=0.5=0.5 to 2 mm/s). The mobile clamp is connected to a slider (EZSM6D040K, Oriental Motor) controlled by a controller unit (ESMC-K2, Oriental Motor) via a wire. The strain rate $U/L$ is in the range 0.025 to 0.2 s-1 in the velocity jump experiment. All experiments are performed at the room temperature $\simeq 25$ C.
@@
Crack propagation — From the images obtained from the digital camera, we track the crack tip position $X$ as a function of time $t$, before the jump, as in Fig. 2 (a), where we showed the results from three samples obtained for the same set of $(U,L)$. The three sets of the results are well overlapped, demonstrating high reproducibility of the slow propagation. As for the fast propagation after the jump, we track the crack position $\Delta X(t)=X(t)-X(t_{c})$ as a function of time $\Delta T=t-t_{c}$ with $t=t_{c}$ corresponding the time of jump, as in Fig. 2 (b), where we show the results from three samples obtained for the same set of $(U,L)$. The three sets of the results are again well overlapped, demonstrating high reproducibility even for the fast propagation.
Figure 3: (a) Velocity $V$ vs. strain $\varepsilon$ for a fixed $L$, exhibiting a collapsed straight region on the log-log scale before the jump and a collapsed point just after the jump. (b) $V$ vs. $\varepsilon$ for a fixed $U$. (c) All the data in (a) and (b) on a rescaled horizontal axis $\varepsilon L$, showing $V$ is a function of $\varepsilon L$ in the straight region and at the points just before and after the jump. (d) $V$ vs. $\varepsilon L$ obtained for various parameters by a different way of painting to clearly observe the crack propagation: partial painting (see the text for the details).
Velocity jump — In Fig. 3 (a), we show the crack-propagation velocity as a function of strain for various pulling speed $U$ at a given sample height $L$ (=15=15 mm), which shows a clear jump. The velocity jump of three orders of magnitude occurs at a critical strain $\varepsilon_{c}\simeq 0.1$ with the velocity just before the jump $V_{B}\simeq 200$ mm/s and that just after the jump $V_{A}\simeq 300$ m/s. The curve shows a clear straight region in the log-log plot before the jump. The straight region and the velocities just before and after the jump, $V_{B}$ and $V_{A}$, are superimposed well even though the data with different $U$ are shown, which demonstrates with a high reproducibility the crack-propagation dynamics leading the jump is independent of the pulling speed $U$.
@@
To see the effect of painting and the initial crack length $a$, we show the data obtained from partially painted films in Fig. 3 (d), with a straight line on the log-log plot with a slope $\alpha\sim 6$. Although we see slight changes in $V_{B}$ and $V_{A}$ as well as in the slope of the straight region from Fig. 3 (c), the basic features are unchanged: (1) $V=F(\varepsilon L)$ except for the small strain region with $F(\varepsilon L)\simeq(\varepsilon L)^{\alpha}$, where the master curve $F(x)$ is universal even if we change $U$, $L$, and $a$ in limited ranges, although the slope $\alpha$ is affected probably by painting. (2) $V_{B}$ and $V_{A}$ are also almost independent from $U$, $L$, $a$, and the way of painting.
@@
Discussion — We find the crack propagating velocity is given as a function of $\Delta L=\varepsilon L$, i.e., the amount of the absolute stretch $\Delta L$, regardless of the sample height $L$. We interpret this unexpected result as follows. There is a characteristic microscopic length scale $\simeq l^{\ast}$ near the crack and practically only the region near the crack tip characterized by $l^{\ast}$ is stretched. In other words, the strain near the crack tip is not governed by $\Delta L/L$ but by $\varepsilon^{\ast}=\Delta L/l^{\ast}$ and the velocity is a function of this effective strain $\varepsilon^{\ast}$. One possible candidate for the scale $l^{\ast}$ may be defined by the characteristic time scale for the glass transition $t_{0}$ ($\simeq$ tens of milliseconds) we obtain from the rheology experiment and the velocity just before the jump $V_{B}\simeq 200$ mm/s: $l^{\ast}=V_{B}t_{0}\simeq$ a few mm. In this respect, the experiment with smaller $L$ with $L<l^{\ast}$ is interesting, although technically challenging.
@@
The elastic modulus and the density of the present sample are $E\simeq 500$ MPa and $\rho\simeq 1000$ kg/m3 at the room temperature. The corresponding velocity of elastic wave is $v\simeq\sqrt{E/\rho}\simeq 500$ m/s, which gives a practical maximum of the crack propagation speed and is comparable with the velocity after the jump $V_{A}\simeq 300$ m/s: After the jump, the velocity reaches the maximum speed. This is in contrast with the case of elastomers, where $V_{A}$ is a few orders of magnitude slower than the velocity of the elastic wave $v\simeq 50$ m/s evaluated for the elastic modulus of the rubbery regime $E\simeq$ a few MPa (e.g., Ref[16]). Because of this, for elastomers we observe a ”fast-mode” in the velocity range from $V_{A}$ to $v$, in which $V$ scales with some power of $\varepsilon$. However, it could be possible that we observe velocity increase after the jump even in the present case if the sample width is much longer, or we perform the static experiment with minimizing the preparation time for relaxation of the set samples.
@@
Conclusion — We found a clear velocity jump in a highly reproducible manner for a polymer film we use on a daily base under the dynamic experiment. The velocity - strain relation can be collapsed onto a master curve after rescaling, which is universal for the stretching speed and the sample size in limited ranges. Our study suggests that the velocity jump can be widely observed in various polymer films if they exhibit a glass transition in rheology measurements, which underscores the potential of the dynamics crack propagation experiment for developing tough polymers.
Acknowledgments — We thank Professor Takashi Uneyama (Nagoya University) for discussion on rheological properties of the sample. We are grateful to Anton Paar Japan K.K. for helpful technical support concerning rheological measurements. This work was supported by JSPS KAKENHI Grant Number JP19H01859 and JP24K00596.
@@
Figure 5: Complex modulus and $\tan\delta$ as a function of temperature at $f=1$ and $10$ Hz.
Appendix: Mechanical Responses
Stress vs. Strain —
In Fig. 4, we show the stress-strain curve measured by two methods, one using a rheometer (MCR 302, Anton Paar) and the other using a hand-made stretching apparatus based on the force gauge and slider system explained in the text. In the former method, we use small samples of width 10 mm and of height $L=18$ mm. In the latter method, we use larger samples of width 150 mm and of height 150 mm. The sample height $L$ determines the strain rate, which is in the range 0.025 to 0.2 s-1 in the velocity jump experiment. In each of plot in Fig. 4 (a) to (d), there are two curves of the same color, which correspond to the data obtained under the same condition but with different sample and thus demonstrate the sample dependence of the results.
@@
In Fig. 4 (b), we compare the results obtained by the two methods. The agreement of the results is good in the region $\varepsilon<\varepsilon_{c}$, but the strain and stress at the breakage is significantly larger for larger samples. This difference may be due to the quality of cutting the sample. The free edge of large samples is cut in the factory at the fabrication in the company, while that of smaller samples is cut in our laboratory. The sharpness of the edge is lower for smaller samples used for the rheometer, which leads to smaller strains at the breakage.
@@
Rheological properties — In Fig. 5, we show results of rheological measurements performed by the rheometer (MCR 302, Anton Paar) at the frequency $f=1$ and 30 Hz. Although PVDC is a crystalline polymer from the monomer structure, the degree of crystallization of the sample may be relatively low and thus transparent. As a result, we see the first relaxation of amorphous structure and the second of crystalline structure in the plot, which respectively correspond to the first and second peak of $\tan\delta$.  Judging from the first peak of $\tan\delta$, the glass transition occurs at $T\simeq$ $13$ C and $f=1$ Hz and at $T\simeq$ $22$ C and $f=30$ Hz. We conclude that the characteristic time scale for the glass transition is $\tau\simeq 1/f\simeq$ tens of milliseconds at the room temperature ($T\simeq$ $25$ C). The wavy feature observed in the plots of $E^{\prime\prime}$ and $\tan\delta$ in the region T=50T=-50 to 15-15 C at $f=1$ Hz is due to insufficient torque and the sharp drop of $^{\prime}, E′′E^{\prime\prime}$, and $\tan\delta$ is due to inertia. These technical artifacts could be ameliorated but, practically, the conclusion on the characteristic time-scale of the glass transition would not be changed.

### v2

@@ Abstract
The velocity jump found in the crack propagation for rubbers has been a powerful tool for developing tough rubber materials. Although it is suggested by a theory that the jump could be observed widely for viscoelastic materials, the report on a clear jump is very limited and, even in such a case, reproducibility is low, except for elastomers. Here, we use a mundane food-wrapping film as a sample and observe the crack propagation velocity with pulling the sample at a constant speed in the direction perpendicular to the crack. As a result, we find the jump occurs at a critical strain with high reproducibility. Remarkably, the plot of the crack-propagation velocity as a function of strain can be collapsed onto a master curve by an appropriate rescaling, where the master curve is found to be universal for change in the pulling speed and in the sample height. The result reveals a key parameter for the jump is the strain, suggesting the existence of a small length that characterizes the deformation along the crack. The present study shows that the jump could be observed if the material exhibits a glass transition in rheology, even if more than one structural relaxation occur as in many viscoelastic materials, although a single relaxation of the glass structure is assumed in the previous theory. This finding suggests the possibility that the velocity jump could be observed in any viscoelastic materials that exhibits a glass transition. Our results set limitations on future theories and open an avenue for the velocity jump to become a tool for developing a wide variety of tough polymer-based materials.
keywords:
Crack propagation , Velocity jump , Polymer sheets , Viscoelasticity , Rubbers and elastomers
journal: J Mech Phys Solids
[…]
Physics Department and Soft Matter Center, Ochanomizu University, 2-1-1
Ohtsuka, Bunkyo-ku, Tokyo 112-8610, Japan
1 Introduction
To develop tough polymer-based materials are crucial in our modern life. Accordingly, the fracture of polymer based materials has been studied intensively [1, 2, 3, 4, 5, 6], but remains an active area of research. One of the principal topics in the study of polymer fracture is the crack propagation, which can be studied through various methods: the cyclic tensile test [7, 8], trouser test [9, 10], static tensile test [11, 12], and so forth [13, 14].
@@ 1 Introduction
In 2017, a theoretical explanation was proposed based on a simple viscoelastic model, which suggests the jump can be observed for a wide range of viscoelastic materials [18]. The jump was reproduced in numerical simulation [17], and the physical picture for the origin of the velocity jump proposed in Ref[18], the glass transition at the crack tip, is strengthened by a study performed under a collaboration among theoretical, numerical and experimental groups [20].
@@
As a matter of fact, the velocity jump was not observed for the porous polypropylene, when the crack propagation was observed under the static boundary condition [24]. However, the jump was observed in a dynamic experiment, in which the crack propagation was observed under a dynamic boundary condition [23]: we first introduce a short crack at one edge of the sample, and then start stretching the sample at a fixed speed to observe the crack propagation [25]. The reason the jump was not observed under the static experiment may be that the sample is relaxed during the preparation for giving a fixed stretch, while such effects of relaxation can be reduced in the dynamic experiment. This higher sensitivity for detecting the velocity jump is one of the advantages of the dynamic experiment over the static one.
@@
Here, we use a mundane food-wrapping film as a sample and observe the crack propagation under the dynamic boundary condition. As a result, we find the jump occurs at a critical strain with high reproducibility. The crack-tip position as a function of time is obtained using two cameras, one for capturing the slow crack propagation before jump and the other with a high frame rate for the fast crack propagation after jump. The plot of the crack-propagation velocity as a function of strain thus obtained shows a little sample dependence: for a fixed pulling velocity and a fixed sample size, we obtained quantitatively similar curves for different samples. Furthermore, we find that the velocity curve is independent of the pulling velocity but dependent on the sample height, while they can be collapsed onto a master curve by rescaling: the master curve thus obtained is universal for change in the pulling speed and the height of the sample. The result reveals a key parameter for the jump is the strain, suggesting the existence of a small length that governs the deformation along the crack. The present study sets limitations on future theories and opens an avenue for the velocity jump to become a tool for developing a wide variety of tough polymer-based materials important for our life. For example, the present study will be relevant for resolving one of urgent issues in modern society, the fire accident of mobile battery by toughening a polymer sheet called the separator in secondary batteries, the deterioration of which leads to the fire accident. Note that the separator is a polymer sheet similar to the sample used in the present study in the two respects: both sheets have similar thicknesses and both are made of crystalline polymers.
2 Materials
@@
From the strain-stress curve obtained at the room temperature at the strain rate $\simeq 0.002$ to 0.07 s-1, the elastic modulus is around 500 MPa and increases with the strain rate. From rheological measurements, the sample exhibits a glass transition at the time scale around tens of milliseconds at the room temperature, where the elastic modulus at the glass phase $\simeq 5$ GPa drops to a smaller value $\simeq 500$ MPa. See Appendix for further details on mechanical and rheological properties.
@@
3 Experimental
As illustrated in Fig. 1, we cut a sheet sample from a roll of width $W$ (=150=150 mm) and set the sheet bottom to the fixed clamp and the sheet top to the mobile clamp so that the sample height at the zero strain becomes $L$ (=10=10 to 25 mm), i.e., the distance between the bottom and top clamps is set to $L$ at the zero strain. We introduce a short line crack of length $a$ at one of the free edge, before we start stretching the mobile clamp at a constant speed $U$ (=0.5=0.5 to 2 mm/s). The mobile clamp is connected to a slider (EZSM6D040K, Oriental Motor) controlled by a controller unit (ESMC-K2, Oriental Motor) via a wire. The strain rate $U/L$ is in the range 0.025 to 0.2 s-1 in the velocity jump experiment. All experiments are performed at the room temperature $\simeq 25$ C.
@@ 3 Experimental
4 Crack propagation
From the images obtained from the digital camera, we track the crack tip position $X$ as a function of time $t$, before the jump, as in Fig. 2 (a), where we showed the results from three samples obtained for the same set of $(U,L)$. The three sets of the results are well overlapped, demonstrating high reproducibility of the slow propagation. As for the fast propagation after jump, we track the crack position $\Delta X(t)=X(t)-X(t_{c})$ as a function of time $\Delta T=t-t_{c}$ with $t=t_{c}$ corresponding the time of jump, as in Fig. 2 (b), where we show the results from three samples obtained for the same set of $(U,L)$. The three sets of the results are again well overlapped, demonstrating high reproducibility even for the fast propagation.
Figure 3: (a) Velocity $V$ vs. strain $\varepsilon$ for a fixed $L$, exhibiting a collapsed straight region on the log-log scale before the jump and a collapsed point just after the jump. (b) $V$ vs. $\varepsilon$ for a fixed $U$. (c) All the data in (a) and (b) on a rescaled horizontal axis $\varepsilon L$, showing $V$ is a function of $\varepsilon L$ in the straight region and at the points just before and after the jump. (d) $V$ vs. $\varepsilon L$ obtained for various parameters by a different way of painting to visualize the crack propagation: partial painting (see the text for the details).
5 Velocity jump
In Fig. 3 (a), we show the crack-propagation velocity as a function of strain for various pulling speed $U$ at a given sample height $L$ (=15=15 mm), which shows a clear jump. The velocity jump of three orders of magnitude occurs at a critical strain $\varepsilon_{c}\simeq 0.1$ with the velocity just before the jump $V_{B}\simeq 200$ mm/s and that just after the jump $V_{A}\simeq 300$ m/s. The curve shows a clear straight region in the log-log plot before the jump. The straight region and the velocities just before and after the jump, $V_{B}$ and $V_{A}$, are superimposed well even though the data with different $U$ are shown, which demonstrates with a high reproducibility the crack-propagation dynamics leading the jump is independent of the pulling speed $U$.
@@ 5 Velocity jump
To see the effect of painting and the initial crack length $a$, we show the data obtained from partially painted films in Fig. 3 (d), with a straight line on the log-log plot with a slope $\alpha\sim 6$. Although we see slight changes in $V_{B}$ and $V_{A}$ as well as in the slope of the straight region from Fig. 3 (c), the basic features are unchanged: (1) $V=F(\varepsilon L)$ except for the small strain region with $F(\varepsilon L)\simeq(\varepsilon L)^{\alpha}$, where the master curve $F(x)$ is universal even if we change $U$, $L$, and $a$ in limited ranges, although the slope $\alpha$ is affected by the way of painting (we changed $a$ in the whole painting experiment, which suggests the result is independent of $a$). (2) $V_{B}$ and $V_{A}$ are also universal, i.e., almost independent of $U$, $L$, $a$, and the way of painting.
@@
6 Discussion
We find the crack propagating velocity is given as a function of $\Delta L=\varepsilon L$, i.e., the amount of the absolute stretch $\Delta L$, regardless of the sample height $L$. We interpret this unexpected result as follows. There is a characteristic microscopic length scale $\simeq l^{\ast}$ near the crack and practically only the region near the crack tip characterized by $l^{\ast}$ is stretched. In other words, the strain near the crack tip is not governed by $\Delta L/L$ but by $\varepsilon^{\ast}=\Delta L/l^{\ast}$ and the velocity is a function of this effective strain $\varepsilon^{\ast}$. One possible candidate for the scale $l^{\ast}$ may be defined by the velocity just before the jump $V_{B}\simeq 200$ mm/s and the characteristic time scale for the glass transition $t_{0}$ ($\simeq$ tens of milliseconds) that we obtain from the rheology experiment: $l^{\ast}=V_{B}t_{0}\simeq$ a few mm. In this respect, the experiment with smaller $L$ with $L<l^{\ast}$ is interesting, although technically challenging.
@@ 6 Discussion
The elastic modulus and the density of the present sample are $E\simeq 500$ MPa and $\rho\simeq 1000$ kg/m3 at the room temperature. The corresponding velocity of elastic wave is $v\simeq\sqrt{E/\rho}\simeq 500$ m/s, which gives a practical maximum of the crack propagation speed and is comparable with the velocity after the jump $V_{A}\simeq 300$ m/s: After the jump, the velocity reaches the maximum speed. This is in contrast with the case of elastomers, where $V_{A}$ is a few orders of magnitude slower than the velocity of the elastic wave $v\simeq 50$ m/s evaluated for the elastic modulus of the rubbery regime $E\simeq$ a few MPa (e.g., Ref[16]). Because of this, for elastomers we observe a ”fast-mode” in the velocity range from $V_{A}$ to $v$, in which $V$ scales with some power of $\varepsilon$. However, it could be possible that we observe velocity increase after the jump even in the present case if the sample width is much longer, or we perform the static experiment with minimizing the preparation time for relaxation in setting samples.
@@
7 Conclusion
We found a clear velocity jump in a highly reproducible manner for a polymer film we use on a daily base under the dynamic experiment. The velocity - strain relation can be collapsed onto a master curve after rescaling, which is universal for the stretching speed and the sample size in ranges we investigated. This suggests the jump is governed by the strain and there exists a small length characterizing the deformation along the crack. The rheological measurement suggests that the velocity jump could be observed not only in polymer films with a single relaxation, which is assumed in the previous theory, but also in films with multi-relaxations, if they exhibit a glass transition. Considering that many viscoelastic materials shows a glass transition with multi-relaxations, the present results underscore the potential of the dynamic crack-propagation experiment for developing tough polymers important for our life such as the separator in secondary batteries and the need for a theory based on viscoelasticity with multi-relaxations.
We thank Takashi Uneyama (Nagoya University) and Naoyuki Sakumich (Tokyo University) for discussions respectively on rheological properties of the sample and on general features of the crack-propagation velocity. We are grateful to Anton Paar Japan K.K. for helpful technical support concerning rheological measurements. This work was supported by JSPS KAKENHI Grant Number JP19H01859 and JP24K00596 and a grant from Eno Science Foundation.
Appendix: Mechanical Responses
@@ 7 Conclusion
Figure 5: Complex modulus and $\tan\delta$ as a function of temperature at $f=1$ and $30$ Hz.
7.1 Stress vs. Strain
In Fig. 4, we show the stress-strain curve measured by two methods, one using a rheometer (MCR 302, Anton Paar) and the other using a hand-made stretching apparatus based on the force gauge and slider system explained in the text. In the former method, we use small samples of width 10 mm and of height $L=18$ mm. In the latter method, we use larger samples of width 150 mm and of height 150 mm. The sample height $L$ affects the strain rate, which is in the range 0.025 to 0.2 s-1 in the velocity jump experiment. In each of plot in Fig. 4 (a) to (d), there are two curves of the same color, which correspond to the data obtained under the same condition but with different sample and thus demonstrate the sample dependence of the results.
@@
In Fig. 4 (b), we compare the results obtained by the two methods. The agreement of the results is good in the region $\varepsilon<\varepsilon_{c}$, but the strain and stress at the breakage is significantly larger for larger samples. This difference may be due to the quality of cutting the sample. The free edge of large samples is cut in the factory at the fabrication in the company, while the edge of smaller samples is cut in our laboratory. The sharpness of the edge is lower for smaller samples used for the rheometer, which leads to smaller strains at the breakage.
@@
7.2 Rheological properties
In Fig. 5, we show results of rheological measurements performed by the rheometer (MCR 302, Anton Paar) at the frequency $f=1$ and 30 Hz. Although PVDC is a crystalline polymer from the monomer structure, the degree of crystallization of the sample may be relatively low and thus transparent. As a result, we see the first relaxation of amorphous structure and the second of crystalline structure in the plot, which respectively correspond to the first and second peak of $\tan\delta$. Judging from the first peak of $\tan\delta$, the glass transition occurs at $T\simeq$ $13$ C and $f=1$ Hz and at $T\simeq$ $22$ C and $f=30$ Hz. We conclude that the characteristic time scale for the glass transition is $\tau\simeq 1/f\simeq$ tens of milliseconds at the room temperature ($T\simeq$ $25$ C). The wavy feature observed in the plots of $E^{\prime\prime}$ and $\tan\delta$ in the region T=50T=-50 to 15-15 C at $f=1$ Hz is due to insufficient torque and the sharp drop of $^{\prime}, E′′E^{\prime\prime}$, and $\tan\delta$ is due to inertia. These technical artifacts could be ameliorated but, practically, the conclusion on the characteristic time-scale of the glass transition would not be changed.